If the effective mass of electron in copper is m

=1.01 m

e

(m

e

m∗=1.01 me(me = mass of an electron) and the electric conductivity of copper is 5.76 ×10

7

 Ω

−1

 m

−1

 with E

F

=7

×107 Ω−1 m−1 with EF=7 eV then calculate the Fermi mean free path for copper.


1
Expert's answer
2021-09-23T08:31:47-0400
"\Lambda=\frac{\gamma\sqrt{2mE}}{q^2n}\\
\Lambda=3.22\cdot10^{-21}m"